# sample.bib

### v1

@@ 3 Analyses
where $\sigma$ is defined in (2) and the expectation is taken with respect to the i.i.d. sampling of $\mathcal{D}\sim\mathcal{Z}$ and the Gaussian noise added at the end of $U$.
Comparison with Prior Work.
To review, our algorithm is a first-order, black-box algorithm that provides $(\epsilon,\delta)$ unlearning while also maintaining performance and computational efficiency. We next compare our algorithm with the prior literature in this area. Earlier in Table 1, we review our algorithmic differences with existing convex and nonconvex certified unlearning algorithms. These algorithms all involve injecting some (Gaussian) noise to render the algorithm outputs probabilistically indistinguishable, whether it is added to the objective function, to the final model weights, or at each step of the training process. We are interested in comparing the noise-unlearning tradeoff of each of these algorithms.
Algorithm Standard Deviation of Injected Gaussian Noise
Newton Step [Guo] $\sigma=\frac{4LG^{2}\sqrt{2\log(1.5/\delta)}}{\lambda^{2}(n-1)\epsilon}$
Descent-to-delete (D2D) [neel21a] $\sigma=\frac{4\sqrt{2}M(\frac{L-\lambda}{L+\lambda})^{K}}{\lambda n(1-(\frac{L-\lambda}{L+\lambda})^{K})(\sqrt{\log(1/\delta)+\epsilon}-\sqrt{\log(1/\delta)})}$
Langevin Unlearning [chien2024langevin] No closed-form solution for $\sigma$ in terms of $\epsilon,\delta$
Constrained Newton Step [zhang2024towards] $\sigma=\frac{\left(\frac{2R(PR+\lambda)}{\lambda+\lambda_{min}}+\frac{32\sqrt{\log(d/\rho)}}{\lambda+\lambda_{min}}+\frac{1}{8}\right)LR\sqrt{2\log(1.25/\delta)}}{\epsilon}$
Hessian-free unlearning [qiao2024efficientgeneralizablecertifiedunlearning] $\sigma=2\eta G\frac{\sqrt{2\log(1.25/\delta)}}{\epsilon}\zeta_{T}^{-U}$
Our Work (R2D) $\sigma=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon}$
Table 2: Comparison of certified unlearning algorithms and their noise guarantees. For the other algorithms referenced in this table, $K$ is the number of unlearning iterates, $\lambda$ is the regularization constant or strongly convex parameter, $M$ is the Lipschitz continuity parameter, $d$ is the parameter dimension. We also have $||\theta||\leq R$ for algorithms that require uniformly bounded model weights. For [zhang2024towards], $P$ represents the Lipschitz constant of the Hessian, $\lambda_{min}$ represents the minimum eigenvalue of the Hessian, and $\rho$ represents a probability less than $1$. For [qiao2024efficientgeneralizablecertifiedunlearning], $\zeta_{T}^{-U}$ is a constant dependent on the upper and lower bounds of the Hessian spectrum that grows with the number of learning iterates $T$ for nonconvex objective functions. See Appendix A for an in-depth discussion of [chien2024langevin].
The last column of Table 2 lists the standard deviation $\sigma$ in terms of $\epsilon$, $\delta$, and other problem parameters for the other certified unlearning algorithms. Our result is analogous to that of [neel21a], due to algorithmic similarities such as weight perturbation at the end of training. However, the required noise in [neel21a] decays exponentially with unlearning iterations, while the noise decreases more slowly with increasing $K$ for our algorithm. This difference can be attributed to the strongly convex setting in [neel21a], which attracts trajectories to a global minimum.
Table 2 highlights some advantages our algorithm has over existing approaches. First, our theoretical guarantees do not require a uniform bound on the model weights $||\theta||\leq R$, nor does $\sigma$ depend on $R$, in contrast to [chien2024langevin] and [zhang2024towards]. Unlearning guarantees with strong dependence on $R$ are problematic, as illustrated in the following example. From the Gaussian mechanism of differential privacy, we have the following theorem.
Theorem 3.5.
@@
[dwork2014algorithmic] Let $X\sim\mathcal{N}(\mu,\sigma^{2}\mathbb{I}_{d})$ and $Y\sim\mathcal{N}(\mu^{\prime},\sigma^{2}\mathbb{I}_{d})$. Suppose $\lVert\mu-\mu^{\prime}\lVert_{2}\leq\Delta$. Then for any $\delta>0$, $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if
@@
Suppose we have an arbitrary learning algorithm that outputs some model weights $\theta$ perturbed by Gaussian noise $\xi\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$ and an arbitrary unlearning algorithm that also outputs some other model weights $\theta^{\prime}$ perturbed by Gaussian noise $\xi^{\prime}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{d})$. Then we have
$\tilde{\theta}=\theta+\xi$
$\tilde{\theta^{\prime}}=\theta^{\prime}+\xi^{\prime}$
$||\theta-\theta^{\prime}||\leq 2R.$
Then by Theorem 3.5, our unlearning algorithm is $(\epsilon,\delta)$ unlearning as long as
$\sigma\geq\frac{\sqrt{2}R}{\sqrt{\log(1/\delta)+\epsilon}-\sqrt{\log(1/\delta)}}.$
In conclusion, an algorithm that outputs arbitrary weights in the feasible set can obtain unlearning with $\sigma=O(R)$, which is of the same order as the dependencies in [chien2024langevin] and [zhang2024towards].
Second, we note that [zhang2024towards] and [qiao2024efficientgeneralizablecertifiedunlearning] require a fixed injection of noise that does not vary with the number of data points or iterations, thereby implicitly upper bounding the utility of the resulting model. In contrast, our result clearly delineates the privacy-utility-efficiency tradeoff of our algorithm. Our noise also scales inversely with the size of the dataset $n$, such that unlearning on large datasets requires less perturbation. Finally, generally speaking, our algorithm requires fewer assumptions than prior works. We do not require Lipschitz continuity of the Hessian [zhang2024towards] or uniformly bounded Hessian eigenvalues [qiao2024efficientgeneralizablecertifiedunlearning] [zhang2024towards].
4 Conclusion
In this work, we introduce a novel certified unlearning algorithm for nonconvex functions that addresses the limitations of prior algorithms. We propose a first-order method that is black-box, in that it can be applied to a model pretrained with gradient descent, and does not have any dependence on the parameter set radius. In general, our algorithm is simple, easy to implement, and requires fewer assumptions than prior work.
A weakness of our algorithm is that it requires the storage of the model checkpoint for later unlearning. In addition, if the checkpoint was not saved during training, the full dataset is required to compute the checkpoint via the proximal point algorithm. However, once the checkpoint is saved, unlearning can occur with access to only the retained dataset.
Future directions of this research include adapting the algorithm to stochastic gradient descent, which is more commonly used for deep neural networks. However, extending the current analysis to the stochastic setting is challenging because for nonconvex functions, the noise at each step causes gradient descent trajectories to diverge exponentially from each other. We will also conduct experiments to empirically compare with other algorithms.
[…]
Appendix A Additional Discussion of Related Work
The work [chien2024langevin] proposes using noisy projected gradient descent to achieve certified unlearning, leveraging the existence and uniqueness of the limiting distribution of the training process as well as the boundedness of the projection set. Their guarantee shows that the privacy loss $\epsilon$ decays exponentially with the number of unlearning iterates.
However, $\sigma$, the noise added at each step, is defined implicitly with no closed-form solution. We can only assert that for a fixed number of iterations $K$, $\sigma$ must be at least $O(R)$ to obtain $\epsilon=O(1)$. Because $\sigma$ cannot be defined explicitly, it is difficult to implement this algorithm for a desired $\epsilon$. For example, when performing experiments for the strongly convex setting, which is a simpler mathematical expression, they require an additional subroutine to find the smallest $\sigma$ that satisfies the target $\epsilon$. As for the nonconvex setting, they state that “the non-convex unlearning bound… currently is not tight enough to be applied in practice due to its exponential dependence on various hyperparameters.”
Appendix B Proof of Theorem 3.1
Proof.
Let $Z$ be a dataset of $m$ data points we would like to unlearn, where $m<n$. Let $\mathcal{D}$ represent the original full dataset and $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$. Without loss of generality, define
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta)$
$f_{\mathcal{D}^{\prime}}(\theta)=\frac{1}{n-m}\sum_{i=1}^{n-m}f_{z_{i}}(\theta)$
@@ Appendix A Additional Discussion of Related Work
$f_{\mathcal{D}}=\frac{n-m}{n}f_{\mathcal{D}^{\prime}}(\theta)+\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta)$
$f_{\mathcal{D^{\prime}}}=\frac{n}{n-m}(f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta))$
@@
$\theta_{0}$ $=\theta^{\prime}_{0}$ (6)
$\theta_{t}$ $=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})$ (7)
$\theta^{\prime}_{t}$ $=\theta^{\prime}_{t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{t-1})$ (8)
@@
$\lVert\theta_{t}-\theta^{\prime}_{t}\rVert\leq\frac{2Gm}{Ln}[(1+\frac{\eta Ln}{n-m})^{t}-1]$
@@
$\nabla f_{\mathcal{D^{\prime}}}(\theta)=\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta))$
$\theta^{\prime}_{t}=\theta^{\prime}_{t-1}-\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}))$
So we have for $\Delta_{t}=\theta_{t}-\theta^{\prime}_{t}$
$\Delta_{t}=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}))$
$=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})$
$=\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})$
$=\Delta_{t-1}+\eta\frac{m}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{n}}(\theta^{\prime}_{t-1})$
We can take the absolute value of each side and by the triangle inequality
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}||\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}||\nabla f_{z_{n}}(\theta^{\prime}_{t-1})||$
Since the gradient on each data point is bounded by $G$, we have
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}G+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}G$
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\frac{2\eta Gm}{n-m}$
@@
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}L||\theta^{\prime}_{t-1}-\theta_{t-1}||+\frac{2\eta Gm}{n-m}$
$||\Delta_{t}||\leq||\Delta_{t-1}||(1+\frac{\eta Ln}{n-m})+\frac{2\eta Gm}{n-m}$
Since we know that $||\Delta_{0}||=0$, evaluating this recursive relationship yields for $t>0$
@@
where the bound on the right hand side decreases monotonically as $K$ increases from 00 to $T$.
Let
$h(K)=((1+\frac{\eta Ln}{n-m})^{T-K}-1)(1+\eta L)^{K}$
@@
$h^{\prime}(K)=(1+\eta L)^{K}[((1+\frac{\eta Ln}{n-m})^{T-K}-1)\log(1+\eta L)-(1+\frac{\eta Ln}{n-m})^{T-K}\log(1+\frac{\eta Ln}{n-m})]$
@@
Therefore, to achieve $\epsilon,\delta$-unlearning, we just need to set
$\sigma=\frac{||\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}||\sqrt{2\log(1.25/\delta)}}{\epsilon}=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon}$
Now we move on to the utility guarantee. For nonconvex smooth functions, we know that for the gradient descent iterates $\theta^{\prime\prime}_{t}$, we have
@@
Let the step-size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have
@@
As before, we first consider the gradient descent iterates on $f_{\mathcal{D}^{\prime}}$. For $\mu$-PL and smooth functions, we know that for the iterates $\theta^{\prime\prime}_{t}$, we have [karimiPL]
@@
Let the step-size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have

### v2

@@ 3 Analyses
where $\sigma$ is defined in (3) and the expectation is taken with respect to i.i.d. sampling of $\mathcal{D}\sim\mathcal{Z}$ and the Gaussian noise added at the end of $U$.
4 Experiments
For more experiment details, including choice of hyperparameters, baseline implementations, hardware, and numerical results, see Appendix D.
(a) Test error vs. rewind percent.
(b) Test error vs. $\epsilon$.
(c) $K$ vs. $\epsilon$.
(d) MIA score vs. $\epsilon$.
(e) Error on unlearned data after unlearning.
Figure 1: Privacy-utility-complexity tradeoff of R2D. The top row shows the experiments on the eICU dataset, and the bottom row shows the experiments on the Lacuna-100 dataset. Figures 1(a), 1(b), 1(c) demonstrate the tradeoff between accuracy and computation/unlearning. Figures 1(d), 1(e) evaluate unlearning as measured by the MIA score and the error on the unlearned data before and after unlearning. The unlearning computations are represented by the rewind percent, or $\frac{K}{T}\times 100\%$.
4.1 Experimental Framework
We test our algorithm (Algorithms 1 and 2) in experimental settings that reflect real-world use cases of unlearning for protecting user privacy. For all experiments, we train a binary classifier with the cross-entropy loss function over a dataset that is naturally split among many different users. To test unlearning, we remove the data associated with a subset of the users ($1$%-$2$% of the data) and observe the impact on the original binary classification task. The unlearned dataset is therefore not i.i.d. to the training or test dataset. Our experimental framework better reflects unlearning in practice, where users may request the removal of their data but they do not each represent a class in the model. This stands in contrast to existing unlearning experiments that unlearn data from a selected class (Guo et al., 2019; Golatkar et al., 2019; Kurmanji et al., 2023), or randomly select samples uniformly from the dataset to unlearn (Zhang et al., 2024; Qiao et al., 2024).  [^7] Code is open-sourced at the following anonymized link: https://github.com/anonymous-1234567/r2d.
4.2 Implementation
We make some approximations of our theory to implement the algorithm in practice. Instead of full-batch gradient descent, we use stochastic gradient descent with a large batch size and sampling without replacement (i.e. shuffling). In addition, while the noise equation (3) assumes constant step size, we use an exponentially decaying step size that allows training to make sufficient progress early on, taking the final step size as $\eta$. Moreover, for $\epsilon>1$, we utilize the bound derived in (Balle & Wang, 2018) to calibrate the noise. For unlearning, we start the step size at the value at the checkpoint. Finally, we estimate the values $L$ and $G$ using sampling approaches outlined in Appendix D. An alternate approach would be to set these parameters as “tunable hyperparameters,” which, as noted in (Zhang et al., 2024) may lead to more imprecise certification guarantees. Future directions of this research include developing theory to incorporate these approximations.
4.3 Datasets
We consider small-scale and large-scale nonconvex settings, performing binary classification on datasets that contain medical or facial information about a set of users.
For the small-scale experiments, we train a multilayer perceptron (MLP) to perform classification on the eICU dataset, a large multi-center intensive care unit (ICU) database consisting of tabular data on ICU admissions (Pollard et al., 2018). Each patient is linked with 1-24 hospital stays. We predict if the length of a hospital stay of a patient is longer or shorter than a week using the intake variables of the Acute Physiology Age Chronic Health Evaluation (APACHE) predictive framework, including blood pressure, body temperature, and age. For unlearning, we unlearn a random subset of patients and their corresponding data.
For the large-scale experiments, we consider the VGGFace2 dataset, which is a dataset of approximately $9,000$ celebrities and their face images from the internet (Cao et al., 2018). We apply the MAAD-Face annotations from (Terhörst et al., 2020) to label each celebrity as male or female, and sample a class balanced dataset of 100 celebrities to form the Lacuna-100 dataset as described in (Golatkar et al., 2019). We train a ResNet-18 neural network model to perform binary gender classification. For unlearning, we remove a random subset of the celebrities and their face images.
4.4 Unlearning Metrics
To empirically evaluate unlearning, we consider the error on the unlearned data before and after unlearning. An increase in error suggests the model is losing information about the data. We also employ black-box membership inference attacks (MIA) that aim to distinguish between unlearned data and data that has never been in the training dataset (Shokri et al., 2016). We train a logistic regression model to classify data samples based on the output (logits and loss) of the unlearned model. We measure the success of the MIA with AUROC on the data samples, computed via repeated $k$-fold cross-validation ($k=100$ for eICU and $k=5$ for Lacuna-100). A higher MIA score corresponds to less successful unlearning. We perform the MIA on the unlearned dataset (one class) and an out-of-distribution (OOD) dataset (the second class) representing data from users absent from the training data. For the Lacuna-100 dataset, we construct an OOD dataset using an additional $100$ users from the VGGFace2 database. For the eICU dataset, we collect data from the test set from users not present in the training set to use in the OOD dataset.
4.5 Results
Figure 1 examines the privacy-utility-complexity tradeoff of our algorithm for $\delta=0.1$ and different values of $\epsilon$, $\sigma$, and $K$. We report the average value over 5 seeds. Figures 1(a) and 1(b) show the accuracy-computation and accuracy-unlearning tradeoff of our algorithm, where “rewind percent” is the percent of training iterations, computed as $\frac{K}{T}\times 100\%$. As expected, increasing the amount of iteration $K$ for the same $\epsilon$ decreases the error, as does decreasing the strength of the unlearning guarantee for the same amount of computation. Figure 1(c) shows the theoretical computation-unlearning tradeoff for fixed noise $\sigma=0.01$ in each setting, for both the numerically computed value of $K$ for values $\epsilon$ and the approximation for $K(\epsilon)$ derived in (3.2).
In Figure 1(d) we empirically assess the extent of unlearning through membership inference attacks. As expected, the MIA score increases as $\epsilon$ increases and the unlearning guarantee is relaxed. For the same $\epsilon$, the MIA is generally more successful when less noise is used. For example, for the Lacuna-100 setting, the MIA is most successful on the noiseless fully retrained model. This suggests that the MIA is picking up on distributional differences between the unlearned data and the OOD data reflected in the output of the model, highlighting both the limitations of relying on the MIA score as an unlearning metric as well as the strong advantages of theoretical $(\delta,\epsilon)$ guarantees. Finally, Figure 1(e) compares the performance of the unlearned model and the original model (without unlearning) on the unlearned data. The increase in error after unlearning suggests that the model is losing information about the samples.
4.6 Comparison with Baseline Methods
Table 2: Comparison of computation time of algorithms for the Lacuna-100 dataset, with $26\%$ or $51\%$ training iterations for R2D. The results for the eICU dataset are in Table 7 in the Appendix.
Algorithm Learning Time Unlearning Time
R2D (26%) 1.47 hours 0.38 hours
R2D (51%) 1.47 hours 0.74 hours
HF 4.23 days 0.00 hours
CNS 0.58 hours 0.32 hours
We compare our algorithm against two other algorithms with theoretical $(\epsilon,\delta)$ unlearning guarantees for nonconvex functions: the Constrained Newton Step method (CNS) (Zhang et al., 2024), a black-box algorithm which involves a single Newton step within a constrained parameter set, and the Hessian-Free method (HF) (Qiao et al., 2024), a white-box algorithm which involves demanding pre-computation and storage of data influence vectors during training, allowing unlearning via simple vector addition later. We do not implement the white-box algorithm (Chien et al., 2024a) because, as stated in their work, “the non-convex unlearning bound… currently is not tight enough to be applied in practice due to its exponential dependence on various hyperparameters.” Although HF technically requires $O(nd)$ storage (and $O(ndT)$ storage during training to compute the unlearning vectors), which is impractically large for our dataset and model sizes, for comparison purposes we implement an $O(md)$ version that only stores the vectors for the data samples we plan to unlearn.
Figure 2: Accuracy and unlearning of baseline unlearning methods. The top row shows the eICU results, and the bottom row shows the Lacuna-100 results. The left plots show the error on the training, test, and unlearned dataset. The right plots show the MIA scores.
Figure 2 shows the performance of the three unlearning algorithms on the training, test, and unlearned dataset as well as the MIA scores after unlearning. In terms of accuracy, R2D generally performs better than CNS and worse than HF, depending on the choice of $K$, due to significant differences in the noise bounds of each algorithm. The MIA results suggest that HF is more successful at defending against membership attacks; however, the low MIA score for CNS is likely due to overfitting of the logistic regression model since the amount of noise added to the model is very high for the range of $\epsilon$ considered.
Table 2 compares the computation time of each approach during learning and unlearning for Lacuna-100, showing that R2D strongly outperforms HF in learning time, and can outperform CNS in the computation saved from unlearning. In contrast, HF, a white-box algorithm, front-loads the computational burden during learning in favor of fast unlearning later. The findings for eICU, in Table 7 in the Appendix, are similar. Ultimately, while both CNS and HF are second-order methods, R2D is a first-order method that provides competitive performance with minimal computation and storage requirements.
5 Conclusion
We propose the first black-box, first-order certified-unlearning algorithm for nonconvex functions, addressing theoretical and practical limitations of prior work. Our algorithm outperforms existing second-order methods in storage, computation, accuracy, and unlearning.
Impact Statement
This paper presents work whose goal is to advance the field of machine unlearning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof of Lemma 2.3
Proof.
We have the following lemma from (Drusvyatskiy, 2017).
Lemma A.1.
(Drusvyatskiy, 2017) If $f(\theta)$ is continuously differentiable with $L$-Lipschitz gradient, $-f(\theta)+\frac{L}{2}||\theta||^{2}$ is convex.
Now we define $\theta^{*}$ as the solution to the proximal problem as follows
$\theta^{*}=prox_{-f,\eta}(\theta_{t+1})=\arg\min_{x}\{-f(x)+\frac{1}{2\eta}||x-\theta_{t+1}||^{2}\},$
which is well-defined due to Lemma A.1 and the fact that $\eta<\frac{1}{L}$. Then the gradient of the objective function is zero at $\theta^{*}$ and thus
$-\nabla f(\theta^{*})+\frac{1}{\eta}(\theta^{*}-\theta_{t+1})=0,$
$\theta^{*}=\theta_{t+1}+\eta\nabla f(\theta^{*}).$
Appendix B Proof of Theorem 3.1
Like prior work in differential privacy and machine unlearning, our work hinges on the Gaussian mechanism for differential privacy, which ensures $(\epsilon,\delta)$-indistinguishability for normal random variables with the same variance.
Theorem B.1.
@@ 5 Conclusion
(Dwork & Roth, 2014) Let $X\sim\mathcal{N}(\mu,\sigma^{2}\mathbb{I}_{d})$ and $Y\sim\mathcal{N}(\mu^{\prime},\sigma^{2}\mathbb{I}_{d})$. Suppose $\lVert\mu-\mu^{\prime}\lVert_{2}\leq\Delta$. Then for any $\delta>0$, $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if
@@
Therefore, to prove Theorem 3.1, we need to bound the distance between the output of the unlearning algorithm and the learning algorithm. We can then add a sufficient amount of noise, scaled by this distance, to achieve $(\epsilon,\delta)$ unlearning.
Proof. Let $Z$ be a dataset of $m$ data points we would like to unlearn, where $m<n$. Let $\mathcal{D}$ represent the original full dataset and $\mathcal{D^{\prime}}=\mathcal{D}\backslash Z$. Without loss of generality, define
$f_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{z_{i}}(\theta),$
$f_{\mathcal{D}^{\prime}}(\theta)=\frac{1}{n-m}\sum_{i=1}^{n-m}f_{z_{i}}(\theta),$
@@
$f_{\mathcal{D}}=\frac{n-m}{n}f_{\mathcal{D}^{\prime}}(\theta)+\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta),$
$f_{\mathcal{D^{\prime}}}=\frac{n}{n-m}(f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}f_{z_{i}}(\theta)).$
@@
$\theta_{0}$ $=\theta^{\prime}_{0},$ (8)
$\theta_{t}$ $=\theta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1}),$ (9)
$\theta^{\prime}_{t}$ $=\theta^{\prime}_{t-1}-\eta\nabla f_{\mathcal{D^{\prime}}}(\theta^{\prime}_{t-1}).$ (10)
@@
$\lVert\theta_{t}-\theta^{\prime}_{t}\rVert\leq\frac{2Gm}{Ln}[(1+\frac{\eta Ln}{n-m})^{t}-1].$
@@
$\nabla f_{\mathcal{D^{\prime}}}(\theta)=\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta)),$
$\theta^{\prime}_{t}=\theta^{\prime}_{t-1}-\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})).$
So we have for $\Delta_{t}=\theta_{t}-\theta^{\prime}_{t}$,
$\Delta_{t}=$ $\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta)-\frac{1}{n}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1})),$
$=$ $\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}),$
$=$ $\Delta_{t-1}-\eta\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}),$
$=$ $\Delta_{t-1}+\eta\frac{m}{n-m}\nabla f_{\mathcal{D}}(\theta_{t-1})+\eta\frac{n}{n-m}(\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1}))-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}\nabla f_{z_{i}}(\theta^{\prime}_{t-1}).$
After taking the absolute value of each side, we obtain by the triangle inequality
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}||\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}||\nabla f_{z_{i}}(\theta^{\prime}_{t-1})||.$
Since the gradient on each data sample is bounded by $G$, we have
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{m}{n-m}G+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||-\eta\frac{1}{n-m}\sum_{i=n-m+1}^{n}G,$
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}||\nabla f_{\mathcal{D}}(\theta^{\prime}_{t-1})-\nabla f_{\mathcal{D}}(\theta_{t-1})||+\frac{2\eta Gm}{n-m}.$
@@
$||\Delta_{t}||\leq||\Delta_{t-1}||+\eta\frac{n}{n-m}L||\theta^{\prime}_{t-1}-\theta_{t-1}||+\frac{2\eta Gm}{n-m},$
$||\Delta_{t}||\leq||\Delta_{t-1}||(1+\frac{\eta Ln}{n-m})+\frac{2\eta Gm}{n-m}.$
Since we have $||\Delta_{0}||=0$, evaluating this recursive relationship yields for $t>0$
@@
where the bound on the right hand side decreases monotonically as $K$ increases from 00 to $T$, as shown by the following. Let
$h(K)=((1+\frac{\eta Ln}{n-m})^{T-K}-1)(1+\eta L)^{K}.$
@@
$h^{\prime}(K)=(1+\eta L)^{K}[((1+\frac{\eta Ln}{n-m})^{T-K}-1)\log(1+\eta L)-(1+\frac{\eta Ln}{n-m})^{T-K}\log(1+\frac{\eta Ln}{n-m})],$
@@
Therefore by Theorem B.1, to achieve $\epsilon,\delta$-unlearning, we need to set the value of $\sigma$ as follows
$\sigma=\frac{||\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}||\sqrt{2\log(1.25/\delta)}}{\epsilon}=\frac{2mG\cdot h(K)\sqrt{2\log(1.25/\delta)}}{Ln\epsilon}.$
Now we prove the utility guarantee. For nonconvex smooth functions, we know by standard analysis that for the gradient descent iterates $\theta^{\prime\prime}_{t}$, we have
@@
Let the step size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have
@@
As before, we first consider the gradient descent iterates on $f_{\mathcal{D}^{\prime}}$. For $\mu$-PL and smooth functions, we know that for the iterates $\theta^{\prime\prime}_{t}$, we have (Karimi et al., 2016)
@@
Let the step size $\eta$ be bounded such that $\eta\leq\frac{n}{2(n-m)L}$, then we have